For the following reaction, 63.2 grams of barium hydroxide are allowed to react with 41.2 grams of sulfuric acid .What is the maximum amount of barium sulfate that can be formed? What is the FORMULA for the limiting reagent?What amount of the excess reagent remains after the reaction is complete? 


1
Expert's answer
2022-01-25T04:01:01-0500

Ba (OH)"_{2(aq)}+" H"_2" SO"_{4(aq)}\to" BaSO"_{4(s)}+2"H"_2O_{(l)}"



Moles of Ba(OH)2

"=\frac{63.2}{171.34}=0.3689"


Moles of H2SO4

"=\frac{41.2}{98.079}=0.4201"


Limiting reagent is Ba(OH)2


Mass of BaSO4 "=0.3689\u00d7233.38"

"=86.09g"


Excess H2SO4 "=0.4201-0.3689"

"=0.0512\>moles"

"=5.0216g"